-15(y-2)-3=-5y-11(3y+7)
-15y-2-3=-5y-33y+7
-15y+5= -38y+7
-53y+12

help!!! Is any of what I did correct?

Answer by nerdybill(7384)

You're not "distributing" correctly...
.
-15(y-2)-3=-5y-11(3y+7)
distributing the -15 to each term inside the parenthesis (left):
-15y+30-3=-5y-11(3y+7)
distributing the -11 to each term inside the parenthesis (right):
-15y+30-3=-5y-33y-77
-15y+27=-28y-77
13y+27=-77
13y=-104
y = -104/13
y = -8